Consider a solution that is 0.10 M in a weak triprotic acid which is represented by the gen- eral formula H3A with ionization constants K1 = 1.0×10−3, K2 = 1.0×10−8, and K3 = 1.0×10−12. What is the pH of the solution?
1
Expert's answer
2012-11-27T08:08:54-0500
K3 is very small,so K2 and K1 is resulting infinding pH
H2A- = H+ + HA-
K2 = [H+]2/H2A-

H2A- = [H+]2/K2

H3A = H+ + [H+]2/K2

K1 = ([H+]2/K2)2/[H3A]

[H+]4 = K1*K22[H3A]

[H+] = (K1*K22[H3A])1/4 = (1.0×10−3*(1.0×10−8)2 * 0.10)1/4 = 1x1010

pH = -log [H+] = 10